# Engagement Maximization

### v4

@@ 4 Implications
Relatedly, our result that the agent’s value function is equal to $\hat{u}(q)$ everywhere implies that a pure diffusion process is infeasible almost everywhere in the continuation region. Because the agent recognizes that the principal will leave her with no surplus, only information that causes her to change her beliefs about the currently optimal action is valuable. If the belief $q_{t}$ is such that one action is strictly optimal given those beliefs (which will be true generically), the principal must offer at least the possibility of jumping to a different region in which another action is optimal; otherwise, the agent will perceive no benefit from the information provided. We interpret this result as suggesting that the principal provides that agent with news articles containing extreme or sensational claims, which should cause the agent to either move her beliefs a lot or not at all, as opposed to providing more nuanced or qualified information.
@@ 5 Extensions
In this section, we consider several extensions to the model analyzed above. We first discuss the possibility that the agent does not fully attend to the principal’s signals, and show that this is never optimal for the agent. Second, we show that commitment to the signal structure plays a minor role in our model. We next consider the case in which the cost of delay is increasing over time, representing either an increasing opportunity cost of time or diminishing utility from using the platform. We then discuss an alternative version of our model in which agents are not capacity constrained and the principal’s objective is to maximize the time until a decision is made. Lastly, we provide an alternative interpretation of our model in the context of a teacher and student.
@@ 5.1 Incentive Compatibility
The first inequality is from the convexity of $\widehat{\mu}$ and $\widehat{\pi}$ being a mean preserving contraction of $q_{\widehat{\tau}}$. The second inequality is the same as the main inequality of Proposition 2. The last equality is from the constraint in Equation 3 being binding for $\pi^{*}$. ∎
@@ 5.2 An Informed Principal
Note that this argument holds even if the agent places the full support of her off-path beliefs on the type with the most reason to deviate (in the spirit of the D1 refinement). This point highlights a key difference between our model and standard Bayesian persuasion models. In standard Bayesian persuasion models, in which the principal cares about the action the agent ultimately chooses, an informed principal whose preferences are aligned with the agent’s has a strong incentive to reveal the true state, which gives refinements like D1 bite. In contrast, in our framework, no principal ever wants to reveal the true state, because each type of principal cares only about maximizing engagement and not about the action the agent ultimately chooses.
@@ 5.3 Increasing Costs of Delay
In this subsection, we consider the case in which the cost of delay is increasing over time. We denote the flow cost of delay at period $t$ by $\kappa(t).$ This case is of interest because it is natural to assume that the utility gained from using a platform declines, and the opportunity cost of time rises, the more time is spent on the platform.
@@
5.3.1 Example (continued)
We illustrate Theorem 2 by continuing the numerical example in Sections 1.1 and 3.5. We maintain the same model primitives except that $\kappa(t)$ no longer constant. The top panel of Figure 4 depicts $\kappa(t)$: $\kappa(t)$ strictly increasing from $[0,2]$ and stays constant afterwards. We choose $\kappa(t)$ take values close to $2$ for the purpose of facilitating comparisons with our previous examples (which had fixed $\bar{\kappa}=2$).
@@ 5.3.1 Example (continued)
The optimal policy is illustrated by the lower panel of Figure 4. The two black curves depict $\mathrm{Supp}(\pi^{*}(\kappa(t)))$. The optimal policy involves the same dilution structure as in Theorem 1, but with posterior beliefs varying over time. If one interpret how extreme the posterior belief is as the decision quality, then decision quality falls over time under increasing time cost.
@@ 5.5 Optimal Policy without Capacity Constraints
This policy is the optimal policy in a special case of our main model. Consider the case of our main model in which $\bar{\kappa}=\chi$ and $G(q)=H(q)=\hat{u}(q)$. In this special case, $\pi^{\max}$ is an optimal policy in our relaxed problem (3), because the constraint in that problem satisfied for any policy, and the process described in Proposition 2 is exactly the process above. The intuition behind this equivalence is that in our main model, it is always optimal for the principal to exhaust the agent’s information processing capacity. If exhausting this capacity necessarily involves satisfying the incentive compatibility constraint, then the principal is free to choose the process that simultaneously exhausts the agent’s capacity and takes as long as possible to reach a decisive belief.
5.6 Engaging Test-Motivated Students
Our model was developed to describe the process of engagement maximization by an internet platform. However, the problem of engagement maximization appears in many contexts. In this subsection, we provide an alternative interpretation of our engagement maximization model in the context of a teacher (the principal) and a student (the agent). We first discuss this interpretation in the context of our general framework, and then provide a concrete example.
Let $R$ be a set of possible responses to a test. Let the states of nature be $X=T\times Q$, where $T$ denotes the true state (a finite set) and $Q$ is the set of possible questions (functions $T\rightarrow R$ describing the correct response conditional on the true state). We will think of the test question realization $\iota\in Q$ as unknown to both the student and teacher (e.g. the test is a standardized test outside the control of the teacher). [^20] A natural extension of this framework would be to consider the case in which the teacher designs and can reveal information about the test questions, in addition to revealing the truth. Exploring this case is beyond the scope of the present paper.
The action $A$ is a function $Q\rightarrow R$ that describes the student’s responses given the question. The student’s utility given action $a\in A$ in state $x=(\omega,\iota)\in T\times Q$ is
$u_{a,x}=v(\sum_{i=1}^{N}\mathbf{1}\{a_{i}(\iota)=\iota_{i}(\omega)\}),$
where $a_{i}(\iota)$ denotes the student’s response to question $i$ and $\iota_{i}(\omega)$ is the correct response to question $i$ given the state $\omega$, and $v(\cdot)$ is an increasing function. That is, the student’s utility is an increasing function of the number of questions answered correctly.
The teacher commits to a teaching strategy without regards to the realization of $\omega$, which determines the sequence of signals the student will receive. [^21] Committing to a teaching strategy before $\omega\in T$ is revealed is analogous, in this context, to the assumption of commitment in Bayesian persuasion models. It avoids the need to consider strategic interactions between teachers with different “types” (knowledge of the truth). The teacher’s objective is to maximize the student’s learning. Let $q_{\omega}=\sum_{\iota\in Q}q_{(\omega,\iota)}$ denote the total probability of the true state $\omega\in T$ under $q\in\mathcal{P}(X)$. We will assume initially that the teacher’s goal is to maximize the expectation of the log of this probability.
The expected utility for the teacher when the teaching strategy is designed (prior to the state being revealed, so that the expectation is taken over stopping beliefs and $\omega$) is
$E^{P}[\ln(q_{\tau,\omega})]=E^{\pi}[\sum_{\omega\in T}q_{\tau,\omega}\ln(q_{\tau,\omega})]$
where $\pi$ is the law of the posterior stopping beliefs $q_{\tau}$ under $P$. The rate at which the student can learn the truth is governed by the mutual information between her prior and posterior over the truth, and it is impossible for the student to acquire any information about the test questions. Define
$H(q)=\begin{cases}\sum_{\omega\in T}q_{\tau,\omega}\ln(q_{\tau,\omega})&\text{proj}(q,\mathcal{P}(Q))=\text{proj}(\bar{q}_{0},\mathcal{P}(Q))\\ \infty&\text{otherwise}.\end{cases}$
This function is infinite when the projections of $q$ and $\bar{q}_{0}$ onto $\mathcal{P}(Q)$ are not equal, ensuring that the student will not acquire information about the test questions until she takes the test, and is otherwise equal to the negative of the Shannon’s entropy of the projection of $q\in\mathcal{P}(X)$ onto $\mathcal{P}(T)$. Under these assumptions, the information processing constraint is exactly that (2) in our main model, and the teacher’s objective is to maximize $E^{P}[H(q_{\tau})]$. Note that we have constructed this example so that the constraint and principal’s objective are governed by the same $H$ function, in keeping with our main example. [^22] This equivalence is a consequence of assuming the log-probability objective and is well-known in the proper scoring rule literature. If we assume that the student has a constant opportunity cost of time spent on the course equal to $\bar{\kappa}$, then this teacher-student model is equivalent to our principal-agent model.
The inherent conflict in this model is that the student is concerned only about the test-relevant portions of the truth, whereas the teacher would like the student to learn both the test-relevant and test-irrelevant aspects of the truth. The teacher must balance covering all aspects of the subject with the need to keep students engaged by providing test-relevant information. Our main results demonstrate that the teacher will drive the student to her reservation value (so that she is indifferent between studying and not) and cause the student to learn more (arrive at more extreme beliefs) than the student would choose for herself.
More subtly, our results show that the information the teacher provides is equivalent to the information the student would choose for herself if she had a lower opportunity cost of time (by Proposition 1 and Theorem 1, under the assumption that $G=H$). One implication of this result, which we highlight in the next example, is that the teacher will never induce the student to acquire test-irrelevant information (i.e. the teacher will “teach to the test”). The teacher optimally prefers to induce the student to acquire more test-relevant information than the student would choose to acquire on her own as opposed to inducing the student to acquire some test-irrelevant information. We illustrate this in the following example. We also show that this conclusion arises as a consequence of the alignment between the teacher’s objective and the student’s information processing constraint (i.e. that $G=H$).
5.6.1 Example (continued)
We adapt the example studied in Sections 1.1 and 3.5 to study engaging test-motivated students. Suppose the true state of the world $T$ consists of both a test-relevant and test-irrelevant dimension, $T=T_{1}\times T_{2}=\{L,R\}\times\{0,1\}$. The student and teacher know that the student will be asked a single question, $Q=\{Q_{0}\}$, whose responses are $R=\{l,r\}$, with $Q_{0}(L0\}=Q_{0}(L1)=l$ and $Q_{0}(R0)=Q_{0}(R1)=r$. That is, the $\{L,R\}$ component of the true state is relevant, and the $\{0,1\}$ component is irrelevant. The student (agent) again faces a binary choice $A=\{l,r\}.$ The student’s utility is one if she answers the question correctly and negative one otherwise, and therefore is identical (ignoring the test-irrelevant dimension) to that in Sections 1.1 and 3.5. The prior belief is uniform. We assume the agent’s information processing capacity is defined by the mutual information of state $i$and the signal, and continue to use the parameters $\bar{\kappa}=2,\chi=1$ and $\rho=1$.
Engagement maximization.
We first study the benchmark case where the teacher maximizes the engagement of the student. This setting is a special case of our main model. By Proposition 1, the teacher’s optimal strategy maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-(2-\lambda)(H(q)-H(\bar{q}_{0}))]$
for some $\lambda<2.$ The student-optimal benchmark maximizes
$\mathbb{E}^{\pi}[\hat{u}(q)-2(H(q)-H(\bar{q}_{0}))].$
Both optimization problem can be written as:
$-\sum_{a,t_{1}}\frac{1}{2}P(a|t_{1})\log(\sum_{t_{1}}\frac{1}{2}P(a|t_{1}))),$
which is equivalent to a static rational inattention problem. [^23] This equivalence is special to mutual information and is the implication of a more general “compression invariance” property introduced by Caplin et al. (2022); Bloedel and Zhong (2020). The analysis above implies that when there is a test-irrelevant state (which is costly to learn about), no information about the state will be acquired either in the student-optimal benchmark or in the principal-agent problem. This is interesting because the student learning about state $t_{2}$ enters the teacher’s payoff function (the measure of engagement), but the teacher optimally provides only test-related information to the student.
Figure 5: Engagement maximization
Figure 5 illustrates the analysis above. The tetrahedron depicts the probability simplex that lies in $\mathbb{R}^{3}$. Each vertex of the tetrahedron is a degenerate belief (labeled by the state that occurs with probability one). The red dot is the uniform prior. The hyperplane represents the cutoff beliefs where the student is indifferent between actions $l$ and $r$. To the right of the hyperplane, $q_{R}>q_{L}$ and $r$ is optimal. Vice versa for the left of the hyperplane. The two blue dots are the student-optimal posterior beliefs. The two green dots are the solution to the principal-agent problem. Evidently, all the dots are one the same straight line, which illustrates that the teacher only releases test-related information to the student, but at a higher precision comparing to what the student prefers.
Knowledge maximization.
Now, we turn to the case where the teacher only cares about how much the student knows about the state $t_{2}\in T_{2}$ (the test-irrelevant dimension). This is a special case of $G\neq H$ in the context of our more general framework. We assume that the teacher’s payoff is $2\mathbb{E^{\pi}}[|q_{t_{2}}-0.5|]$ if the measure of stopping belief is $\pi$, where $q_{t_{2}}$ denotes the probability of $t_{2}=1$ under the student’s stopping belief. Note that this payoff function is equivalent to the instrumental value of information in a hypothetical binary decision problem where the utility of matching (mismatching) the state $y$ is 1 (-$1$). Proposition 1 can be extended to show that the teacher’s optimal policy maximizes the following auxiliary problem (for some $\lambda>0$):
Per Matêjka et al. (2015), the solution is logit, given by
is identical to the student-optimal solution (and independent to $\lambda$). The unknown parameter $\lambda$ can be pinned down by setting the student’s IC condition binding. The analysis above suggests that when the teacher only cares about the test-irrelevant knowledge, she will give the student exactly his preferred test-relevant information and extra knowledge such that the student gets barely enough welfare to be willing to participate.
Figure 6: Knowledge maximization
@@ 6 Conclusion
We have considered the problem of a principal who provides information to an agent so as to maximize the attention the agent pays to the principal’s information (engagement). The agent values this information for instrumental purposes, is rational and Bayesian, and faces a constraint on the rate at which she can process information. Our main results are (i) that by maximizing engagement, the principal leaves the agent no better off than if she could not receive any information at all, and (ii)that the agent will end up holding extreme beliefs, relative to a benchmark in which the agent could choose the information for herself. Our results highlight the pitfalls of presenting users with information for the purposes of maximizing engagement, a standard practice on popular internet platforms.
Appendix A Additional Proofs

### v5

@@ 4.2 Implications of credibility
When beliefs are suspensive but not strictly suspensive, the agent’s value function is equal to $\hat{u}(q_{t})$ at all times. This implies that a pure diffusion process is infeasible almost everywhere in the continuation region. Because the agent recognizes that the principal will leave her with no surplus, only information that causes her to change her beliefs about the currently optimal action is valuable. If the belief $q_{t}$ is such that one action is strictly optimal given those beliefs (which will be true generically), the principal must offer at least the possibility of jumping to a different region in which another action is optimal; otherwise, the agent will perceive no benefit from the information provided. In the internet platform context, we interpret this result as suggesting that the principal provides the agent with news articles containing extreme or sensational claims, which should cause the agent to either move her beliefs a lot or not at all, as opposed to providing more nuanced or qualified information. In the teacher-student context, to engage a test-motivated student, the teacher must provide signals that have some possibility of changing the student’s belief about the correct answer on the test.
4.3 Illustrative Examples
Example 1 (continued).
We illustrate Propositions 6 and 7 by continuing example 1. First, focus on the $\bar{q}_{0}=\frac{1}{2}$ case. As is illustrated in Figure 1, the unique optimal $\pi^{*}$ involves two posterior beliefs $\left\{q^{1},q^{2}\right\}$, where $q^{1}<\underline{q}$ and $q^{2}>\bar{q}$. Per Proposition 2, the dilution of $\pi^{*}$ implements an optimal policy. In this case, the dilution of $\pi^{*}$ is the unique optimal policy. It is apparent from the symmetry of the problem that the value of information acquisition, $\overline{V}_{R}(q,\pi^{*})$, is maximized at $q=\frac{1}{2}$. As a result, the set of suspensive beliefs is a singleton, $\{\bar{q}_{0}\}$. By Proposition 6, beliefs will remain at $\bar{q}_{0}$ until they jump to either $q^{1}$ or $q^{2}$.
When $\bar{q}_{0}<\frac{1}{2}$, the optimal policy is not unique. In Figure 7, we plot the simplest dilution policy, where belief stays at $q_{0}$ until it jumps to the two optimal posteriors $q^{1}$ and $q^{2}.$ However, there can be other optimal policies. In this case, the set of suspensive beliefs is $[q_{0},1-q_{0}]$. Then, a full-commitment policy is optimal as long as beliefs remain in this interval $[q_{0},1-q_{0}]$ until eventually jumping to either $q^{1}$ or $q^{2}$. Specifically, Figure 7 plots one of such policies: when belief is at $q_{0}$ or $1-q_{0}$, it either jumps to the closer boundary, or drifts into the interior of $(q_{0},1-q_{0})$. When belief is interior, it follows a diffusion.
[…]
Figure 5: Dilution policy
[…]
Figure 6: Jump/diffusion policy
[…]
Figure 7: Credible policy
However, if the principal would like the policy to be credible, the interim belief cannot be strictly suspensive. In this example, $\bar{V}_{R}=\underline{V}_{R}>0$ for all $q_{t}\in(q_{0},1-q_{0})$, which is to say that these beliefs are strictly suspensive. It follows that a credible policy must satisfy $q_{t}\in\{q_{0},1-q_{0}\}$. Therefore, a credible belief process can only jump between $q_{0}$ and $1-q_{0}$ or jump to $q^{1}$ and $q^{2}$, as illustrated by Figure 7 (note that the dilution policy is also credible, but the jump/diffusion policy is not credible). Figure 8 illustrates the sample paths of one credible policy, in which beliefs jump between $\bar{q}_{0}$ and $1-\bar{q}_{0}$ before eventually jumping to $q^{1}$ or $q^{2}$.
Figure 8: Sample paths of a credible policy
Example 2 (continued).
We next illustrate the implications of Propositions 6 and 7 by continuing example 2, which discusses our two leading cases.
In the $H=G$ leading case, as discussed above, beliefs move on a line segment. As a result, even though the state space in example 2 is two-dimensional, our analysis for example 1 still applies. Because the prior is uniform, the only suspensive belief is the prior, and consequently the dilution policy is the unique optimal policy. This is illustrated below in Figure 10.
In the decision-irrelevant $G$ leading case, in contrast, there are four possible optimal stopping beliefs under the unique $\pi^{*}$ that solves (R). Because of the symmetry of the problem, we will still have $\bar{V}_{R}=\underline{V}_{R}$, and the set of beliefs that are suspensive but not strictly suspensive are characterized by the level set $\bar{V}_{R}(q,\pi^{*})=0$. These are illustrated by the curves shown in Figure 10.
Any credible, incentive-compatible policy must maintain interim beliefs on these curves. Because they are curves and not straight lines, this is not compatible with beliefs following a diffusion process. But is compatible with a wide range of jump processes, which are the multi-dimensional analogs of the credible policies illustrated in Figures 7 and 8.
Credible interim beliefsR0R1L0L1
Figure 9: Credible beliefs when $H=G$
R0R1L0L1Credible interim beliefs
Figure 10: Credible beliefs with a decision-irrelevant $G$
@@ 5 Extensions
In this section, we consider several extensions to the model analyzed above. We first discuss the possibility that the agent does not fully attend to the principal’s signals, and show that this is never optimal for the agent. Second, we show that commitment to the signal structure plays a minor role in our model (in contrast to the importance of this assumption in the Bayesian persuasion literature). We next consider the case in which the cost of delay is increasing over time, representing either an increasing opportunity cost of time or diminishing utility from using the platform. We then discuss an alternative version of our model in which agents are not capacity constrained and the principal’s objective is to maximize the time until a decision is made.
@@ 5.1 Incentive Compatibility
The first inequality is from the convexity of $\widehat{\mu}$ and $\widehat{\pi}$ being a mean preserving contraction of $q_{\widehat{\tau}}$. The second inequality is the same as the main inequality of Proposition 2. The last equality is from the constraint in Equation R being binding for $\pi^{*}$. ∎
@@ 5.2 An Informed Principal
Note that this argument holds even if the agent places the full support of her off-path beliefs on the type with the most reason to deviate (in the spirit of the D1 refinement). This point highlights a key difference between our model and standard Bayesian persuasion models. In standard Bayesian persuasion models, in which the principal cares about the action the agent ultimately chooses, an informed principal whose preferences are aligned with the agent has a strong incentive to reveal the true state, which gives refinements like D1 bite. In contrast, in our framework, no principal ever wants to reveal the true state, because each type of principal cares only about maximizing engagement and not about the action the agent ultimately chooses.
@@ 5.3 Increasing Costs of Delay
In this subsection, we consider the case in which the cost of delay is increasing over time. We denote the flow cost of delay at period $t$ by $\kappa(t).$ This case is of interest because it is natural to assume that the utility gained from using a platform declines, and the opportunity cost of time rises, the more time is spent on the platform. The same can be said for the utility of learning and the opportunity cost of time in the teacher-student context.
@@
5.3.1 Example 1 (continued).
We illustrate Theorem 3 by continuing our example 1. We maintain the same model primitives except that $\kappa(t)$ no longer constant. The top panel of Figure 11 depicts $\kappa(t)$: $\kappa(t)$ strictly increasing from $[0,2]$ and stays constant afterwards. We choose $\kappa(t)$ take values close to $2$ for the purpose of facilitating comparisons with our previous examples (which had fixed $\bar{\kappa}=2$).
@@
The optimal policy is illustrated by the lower panel of Figure 11. The two black curves depict $\mathrm{Supp}(\pi^{*}(\kappa(t)))$. The optimal policy involves the same dilution structure as in Theorem 1, but with posterior beliefs varying over time. If one interprets how extreme the posterior beliefs are as the decision quality, then decision quality falls over time under increasing time costs.
@@ 5.5 Optimal Policy without Capacity Constraints
This policy is the optimal policy in a special case of our main model. Consider the case of our main model in which $\bar{\kappa}=\chi$ and $G(q)=H(q)=\hat{u}(q)$. In this special case, $\pi^{\max}$ is an optimal policy in our relaxed problem (R), because the constraint in that problem satisfied for any policy, and the process described in Proposition 2 is exactly the process above. The intuition behind this equivalence is that in our main model, it is always optimal for the principal to exhaust the agent’s information processing capacity. If exhausting this capacity necessarily involves satisfying the incentive compatibility constraint, then the principal is free to choose the process that simultaneously exhausts the agent’s capacity and takes as long as possible to reach a decisive belief.
@@ 6 Conclusion
We have considered the problem of a principal who provides information to an agent so as to maximize the attention the agent pays to the principal’s information (engagement). The agent values this information for instrumental purposes, is rational and Bayesian, and faces a constraint on the rate at which she can process information. Our main results are (i) that by maximizing engagement, the principal leaves the agent no better off than if she could not receive any information at all, and (ii) that the agent will end up holding extreme beliefs, relative to a benchmark in which the agent could choose the information for herself, and (iii) that optimal signals in the absence of commitment cannot either increase or decrease the value of future information prior to stopping (i.e. that they are suspensive but not strictly suspensive). Our results highlight the distortions created when an information provided maximizes engagement.
Appendix A Omitted Proofs for Section 3
@@ A.1 Proof of Proposition 1
Proof.
Assumption 1 implies that the constraint in (R) must be binding, as otherwise, the optimal $\pi^{*}$ will be fully revealing, leaving the agent with negative surplus. Then, (R) is a convex optimization problem that satisfies the conditions in Theorem 4 of Zhong [2018], which immediately the existence of $\lambda$ and a finite support solution $\pi^{*}$.
Next, we prove $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$. Suppose for the purpose of contradiction that $\mathbb{E}^{\pi^{*}}[J(q)]>0$. For each $q\in\text{Supp}(\pi^{*})$, let $\pi_{q}$ be the probability measure that implements $\bar{J}(q)$. Let $\pi^{\prime}=\mathbb{E}^{\pi^{*}(q)}[\pi^{q}]$. Then,
$\mathbb{E}^{\pi^{\prime}}[G(q)-G(\bar{q}_{0})]=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\mathbb{E}^{\pi_{q_{1}}}[G(q)]-G(q_{1})+G(q_{1})-G(\bar{q}_{0})]$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\mathbb{E}^{\pi_{q_{1}}}[G(q)-G(q_{1})]]+\bar{J}(\bar{q}_{0})$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}[\bar{J}(q_{1})]+\bar{J}(\bar{q}_{0})$
$>$ $\bar{J}(\bar{q}_{0}).$
Meanwhile,
$\mathbb{E}^{\pi^{\prime}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi^{*}(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)\right]-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi^{*}(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{1})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{1})\right]\right]$
$+\mathbb{E}^{\pi^{*}(q_{1})}\left[\hat{u}(q_{1})-\frac{\bar{\kappa}}{\chi}H(q_{1})-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$\geq$ $0$
Therefore, $\pi^{\prime}$ is feasible and strictly improves $\pi^{*}$, contradiction. ∎
A.2 Proof of Proposition 2
Proof.
Take any admissible stopping time $\tau^{\prime}$,
$=$ $\mathbb{E}^{P}[\widehat{\mu}(q)-\bar{\kappa}\tau|\mathcal{F}_{0}].$
The first equality is from the definition of conditional expectations and the process defining $q_{t}$, (5). The first inequality is from the constraint $\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$ in the relaxed problem (R) and $E[\tau|\tau^{\prime}\geq\tau]\leq E[\tau^{\prime}|\tau^{\prime}\geq\tau]$. The second equality is from the memorylessness property of $\tau$, [^24] Specifically, $\mathbb{E}^{P}\left[\tau|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau-\tau^{\prime}|\tau^{\prime}<\tau\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]=\mathbb{E}^{P}\left[\tau|\mathcal{F}_{0}\right]+\mathbb{E}^{P}\left[\tau^{\prime}|\tau^{\prime}<\tau\right]$. and the last from the definition of conditional expectations. Note that if the constraint in (R) is slack, the first inequality is strict if $\tau^{\prime}$ is not equal to $\tau$ $P$-a.e. In this case, the $\alpha$-dilution of $\pi$ strongly implements principal’s utility level $\rho\mathbb{E}^{\pi}[G(q_{t})-G(\bar{q}_{0})]$. ∎
A.3 Proof of 2
Proof.
If $V^{B}(q)>\hat{u}(q)$, then let $\pi$ implement $V^{B}(q)$. Let $\pi^{\prime}$ satisfies $\mathbb{E}^{\pi}[H(q^{\prime})-H(q)]=\mathbb{E}^{\pi^{\prime}}[H(q^{\prime})-H(q)]$ and $\mathbb{E}^{\pi^{\prime}}[G(q^{\prime})-G(q)]>0$. Then, the dilution of a convex combination of $\pi$ and $\pi^{\prime}$ is feasible, incentive compatible for sufficiently small weight on $\pi^{\prime}$ and strictly profitable for the principal. Therefore, $J(q)>0$.
Next, suppose for the purpose of contradiction that $\pi$ implies $V^{B}(\bar{q}_{0})$ but $\mathbb{E}^{\pi}[V^{B}(q)-\hat{\mu}(q)]>0$. For each $q\in\text{Supp}(\pi)$, let $\pi_{q}$ be the probability measure that implements $V^{B}(q)$. Let $\pi^{\prime}=\mathbb{E}^{\pi(q)}[\pi^{q}]$.
$\mathbb{E}^{\pi^{\prime}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]=$ $\mathbb{E}^{\pi(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)\right]-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$=$ $\mathbb{E}^{\pi(q_{1})}\left[\mathbb{E}^{\pi_{q_{1}}}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)-\hat{u}(\bar{q}_{1})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{1})\right]\right]$
$+\mathbb{E}^{\pi(q_{1})}\left[\hat{u}(q_{1})-\frac{\bar{\kappa}}{\chi}H(q_{1})-\hat{u}(\bar{q}_{0})+\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})\right]$
$>$ $V^{B}(\bar{q}_{0})-\hat{u}(\bar{q}_{0}).$
This contradicts the optimality of $\pi$. ∎
A.4 Proof of Proposition 3